You work in a research lab with a chemist who asks you to make 500.0 mL of a formic acid/sodium formate buffer solution with pH = 4.10. Formic acid is HCOOH(aq); sodium formate is NaHCOO(s) and is soluble in water. The total concentration of formic acid + formate ion in the buffer is to be 0.120 M. Ka(HCOOH) = 1.77 × 10–4 (pKa = 3.75).

In the storeroom, there are the following reagents: (i) solid sodium formate, (ii) a 3.00 M HCl(aq) stock solution, (iii) a 3.00 M NaOH(aq) stock solution, and (iv) deionized water. To prepare the prescribed buffer,

(i) (2 pts) how many grams of solid sodium formate are needed?

(ii) (4 pts) how many milliliters of either 3.00 M HCl(aq) or 3.00 M NaOH(aq) are needed? Explain your choice of reagent by writing an appropriate chemical reaction.

(iii) (2 pts) approximately how many milliliters of deionized water are needed?
Answer #1

a) The mass of sodium format required is calculated:

m Format = M * V * MM = 0.12 M * 0.5 L * 68 g / mol = 4.08 g

b) The molar relationship between the components of the buffer is calculated:

n Salt / n Acid = 10 ^ (pH - pKa) = 10 ^ (4.10 - 3.75) = 2.2

It has:

1) n Salt - 2.2 * n Acid = 0

2) n Salt + n Acid = M * V = 0.12 * 0.5 = 0.06

System of equations is applied and you have:

n Salt = 0.04 mol

n Acid = 0.02 mol

The HCl reacts with the acid form and form, the volume of HCl required is calculated:

V HCl = n * 1000 / M = 0.02 * 1000/3 = 6.7 mL

c) The volume of water required is calculated:

V H2O = 500 mL - 6.7 mL = 493.3 mL

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.